OpenStudy (anonymous):

If f(x)=sin(2x) then f(π/4) =

OpenStudy (tkhunny):

Have you considered substituting the value pi/4 into the function definition?

OpenStudy (anonymous):

f(x)=sin(2x) Substituting x=π/4, we find \[f(π/4) =\sin(2\times \frac{\pi}{4})=\sin( \frac{\pi}{2})=1\] Thus \[f(π/4) =1\]

OpenStudy (tkhunny):

{sigh} I guess we'll never know...

OpenStudy (anonymous):

how does sin(π/2)=1

OpenStudy (anonymous):

Watch the trigonometric table...

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

trigonometric table contains the values for all trigonometric functions Visit here http://www.sosmath.com/tables/trigtable/trigtable.html